I am not good with inequalities so i need help with the following...
21-y=(-87)+2y
6-y=(-y)+2
4/3u-6=7/3u+8
7+6z=8z-13
12p=6-3p
5/6t+4=2-1/6t
3+2(k+1)=6+3k
6r+5=8(r+2)-2r
(-2)(2c-4=1/3(-12c+24)
u/0.3=4u+6.28
0.3x-15=0.2x-5

Answer by Bb(19)

ok i'm not going to answer all of these but i'll teach u how to do the first one 21-y=-87+2y ok 1st u'd +87 to 21 and u get 108-y=2y then u add y to 2y and u get 108=3y then divide both sides by 3 and u get 36 do the rest of the questions that way and u'll be fine


i hope this helped u!!